19 solve for x: -5x/7=-10/21
Answer by checkley75(3666)

-5X/7=-10/21 CROSS MULTIPLYING WE GET
-5X*21=-70
-105X=-70
X=-70/-105
X=2/3
PROOF
-5(2/3)/7=-10/21
(-10/3)/7=-10/21
(-10/3)(1/7)=-10/21
-10/21=-10/21